SpaceRef

SpaceRef


ISS On-Orbit Status 11 Mar 2003

Status Report From: NASA HQ
Posted: Tuesday, March 11, 2003

image All ISS systems continue to function nominally, except as noted previously or below.

Before breakfast, all crewmembers completed another one round of the monthly Russian medical experiment PZEh-MO-8 (body mass measurement, BMM), after FE-1 Nikolai Budarin set up the MO-8 "scales" equipment (IMT).  [MO-8 determines the inertial forces caused by the oscillatory motion of a mass driven by two helical metering springs with known spring constant.  By measuring the time period of each oscillation of the unknown mass (the crewmember) and comparing it to the period of a known mass, the crewmember’s mass is calculated by the computer and displayed.]

Later in the day, before breaking down the IMT and stowing it away, Budarin also determined the mass of 14 CO2 absorbent cartridges (PP) by fastening each cartridge to his back before mounting the BMM device and logging the difference to his own without-cartridge mass.  [The 14 PPs in the Service Module (SM) have a guaranteed shelf life through January 2003.  In order to extend the shelf life, their absorption capacity must be ascertained.  Increased mass would indicated reduced capacity.]

Along with MO-8, the crew also completed the monthly PZEh-MO-7 (calf volume measurement) assessment.  [Calf measurements (left leg only) are taken with the ISOG device, a custom-sewn fabric cuff that fits over the calf, using the knee and lower foot as fixed reference points, to provide a rough index of deconditioning in zero-G and effectiveness of countermeasures.]

FE-2/SO Don Pettit continued to work with the Payload Operations Center (POC) in troubleshooting the Microgravity Science Glovebox (MSG).  The troubleshooting is being accelerated by modifying Wednesday's and Friday's crew timeline to allow more power path runs. [After yesterday's re-installation of the P105 and P103 connectors, early analysis has shown again no off-nominal current draw on ESEM3 (exchangeable standard electronic module 3).  Today Pettit is to reconnect E-Box connector P104 and, if nominal, connector P110.  As of now, eight of 13 power connectors have been remated without locating the fault.  In the event that the remaining re-connections also remain fruitless, the possibility of additional fault isolation steps will be worked.  If the problem is still not found after that, the fault tree leads to the conclusion of "nominal configuration", and normal MSG operations would then be requested next week.  Also, before the MSG back panel is finally closed, the failed left filter-bank fan would be replaced with an onboard spare.]

In the continuing interest of conserving precious onboard resources and reducing upmass required on future Progress flights, CDR Ken Bowersox today conducted an inventory audit of crew clothing.  [The audit was broken into several parts: counting unopened bags (contents are known to the ground) and then reporting the contents of any opened or repacked bags, by filling out tables uplinked overnight.]

Nikolai Budarin performed his daily status check on the Russian BIO-5 Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment in the SM.

Later, he completed the daily routine maintenance of the SM's SOSh life support system, including ASU toilet facilities, while Pettit prepared the daily "delta" file of the IMS (inventory management system) database.

At 11:15 am EST, via ARISS "telebridge" station in Kingston, Australia, the Science Officer fielded questions during an amateur radio exchange with Grade 3-5 students at Eugene Field School in Park Ridge, Illinois, with which he already had a very successful ham exchange back in January (1/23).  [More than 600 students attend this school, which counts as its most famous alumni U.S. Senator Hilary Clinton and actor Harrison Ford.]

Later, at 12:54 pm, the crew participated in an educational Q&A downlink for the NASA Langley Research Center, to be used in an upcoming segment of the NASA Science Files television show, a distance learning initiative airing on more than 80 PBS stations and seen by more than three million public schools students.

All crewmembers performed their daily physical exercise program on RED, TVIS and VELO with load trainer.  Subsequently, after his run on the treadmill, Budarin performed the weekly TVIS maintenance.

The Russian/French experiment LSO, using Molniya-SM equipment, continues to operate at SM window #3, looking for thunderstorm clouds and the very rare glow phenomenon of a "sprite".

Don Pettit was thanked by the ground for successfully troubleshooting BC-1 (battery charger #1) of the BSA (battery stowage assembly) yesterday after initiating recharge of EMU batteries.  [BC-1 has a history of shutting down unexpectedly, due to a known software timing issue.  It will eventually be swapped out with a new unit with upgraded software.]

Last night's survey of the portside SM solar array (#2) with the SSRMS (space station remote manipulator system), in particular the Kurs-P feeder cable on the array's reverse side, was completed successfully. [Only half of the coverage could be viewed live at MCC-H due to a ground systems "hiccup", but the full video record was transmitted to TsUP/MCC-Moscow early this morning, to a grateful "Spasibo bolshoe!"]

Yesterday's Mission Configuration Uplink test of the new R3 GNC (guidance, navigation & control) MDM "stale RGA data" software patch was successfully completed.  Its upload is scheduled for tomorrow (3/12), in conjunction with the handover of ISS attitude control (AC) to the Russian segment (RS) at 2:00pm for the first Progress thruster test reboost at 5:23pm, and the return of AC from RS to USOS (U.S. segment) thereafter.  The mini-reboost is planned to produce a delta-V of 1.38 m/sec.  The second Progress thruster manifold will be tested on 3/13, at 6:38pm, for 0.31 m/sec.

The GNC software upload will also include the new PPL (pre-positioned load) file with larger rate divergence threshold values as an interim solution to the RGA (rate gyro assembly) dilemma issue observed last week (see 3/8 Status Report) and still "latched" in the C&W (caution & warning) system.

The Russian software upgrade to the 7.01 version is on standby, pending the conclusion of the U.S. R3 GNC step-up.  Start date for Russian code transition is still 3/17 (next Monday).

ISS flight attitude, currently in earth-"fixed" LVLH (local vertical/local horizontal), will be maneuvered back to XPOP (x-axis perpendicular to orbit plane) Thursday night, immediately after the second Progress thruster test.

The Stage EVA by Sox and Don has been shifted from 3/25 to early April (tentative date: 4/2 or 4/4).  [For acceptable communications coverage, the EVA needs LVLH attitude, which would have required maneuvering from XPOP and back.  Since ISS will go to LVLH anyway in early April, delaying the EVA to that period conserves valuable attitude maneuver propellant.  Preferred date would have been 4/3, but MCC-Moscow requested that day for its periodic solar array efficiency test (one already had to be skipped), for which 4/3 apparently offers rare optimum conditions of near-zero Beta angle and XPOP.  Further study is underway.]

Also in the continuing interest of conserving precious onboard propellants, the U.S. P6 solar array wings 2B and 4B continue to "fly" in dual-angle directed position such that they are feathered in "night glider" configuration, in order to reduce aerodynamic drag during the eclipse portion of the orbit.  To date, this has worked fine, but it remains to be seen whether it contributes noticeably to propellant conservation.

During yesterday's troubleshooting of the IMV (intra-module ventilation), Bowersox discovered in the FGB a collapsed section of fabric air duct, which, when stretched out, oscillates in the air stream and causes the noise reported by the crew.  A video of the duct, which Sox temporarily restrained with a rope, was provided to Moscow for analysis.  [Since the recent IMV fan cleaning in the Lab and Node, ventilation airstream has been stronger in the USOS than in the RS, which probably caused the duct oscillation.]

Carbon dioxide partial pressure (ppCO2) in the Lab was 3.9 mmHg this morning.  Vozdukh continues to run nominally in manual mode.  CDRA activation has not been required.

A hard-disk drive (HDD) failed last week in the U.S. PCS (portable computer system) laptop in the SM and was replaced.  Also, two RPCM (remote power controller module) "health" flags popped up in the USOS.  One is already refreshed, the other was being looked at this morning.

MCC-H Flight Surgeon requested OCA comm. downlink of accumulated MEC (medical equipment computer) files for CEVIS (cycle ergometer with vibration isolation), HRM (heart rate monitor), RED, TVIS, and FFQ (food frequency questionnaire) nutrition log.   In the interest of conserving precious resources, Budarin completed another extensive 3-hr. IMS inventory audit in the FGB "Zarya" module, of up to 40 items shown on the IMS database as stowed behind Panel 227.  Audit objective was to update the status of dust collector/filters and cartridges as well as U.S. hardware availability.

Today's targets for the CEO (crew earth observations program) were Angolan Biomass Burning(this was a well-illuminated pass in fair weather over southwestern and southern Angola.  Although no active burning was anticipated, the crew was to take advantage of nadir views to map burn scar patterns for previous seasons), Industrialized Southeastern Africa (warm, stable high pressure grips southern Africa.  Looking of industrial aerosol build-up in oblique views either side of track), Recife, Brazil (ISS had a nadir pass over this easternmost port city of Brazil), Havana, Cuba (this was a very good pass for the Cuban capital.  As ISS approached the coast from the NW, the crew was to look just right of track for details of the city and its port facilities), Guadalajara, Mexico (nadir pass over Mexico's second largest city. Besides the urban area of the city, crew was to try to document the heavy resort development around Lake Chapala just southeast of Guadalajara), and Tuamotu Archipelago (resuming detailed ISS/CEO documentation of this archipelago.  This pass, in high illumination, tracked over the denser, western portion of this group.  Crew was to use the long lenses to capture details of the coral reef structures in nadir views). CEO images can be viewed at the websites http://eol.jsc.nasa.gov and http://earthobservatory.nasa.gov

ISS Orbit  (as of this morning, 7:03am EST [= epoch]):

  • Mean altitude -- 388.9 km
  • Apogee -- 396.8 km
  • Perigee -- 380.9 km
  • Period -- 92.33 min.
  • Inclination (to Equator) --  51.63 deg
  • Eccentricity -- 0.00118
  • Orbits per 24-hr. day -- 15.59
  • Mean altitude loss in last 24 hours -- 190m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  -- 24576
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see
  •   http://www.hq.nasa.gov/osf/station/viewing/issvis.html

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.